OpenStudy (cherrilyn):

Evaluate the integral using trig substitution. Refer to trig. integrals. . Integral of dx/x^2 times sqrrt of x^2-25.

OpenStudy (cherrilyn):

\[\int\limits_{}^{}dx / x ^{2} \sqrt{x ^{2}-25}\]

OpenStudy (anonymous):

Where is this x^2 in denominator or up top?

OpenStudy (cherrilyn):

it's in the denominator

OpenStudy (anonymous):

I got it, let x= 5 sec theta

OpenStudy (anonymous):

u=(x^2) - 25 = 5 tan theta

OpenStudy (anonymous):

do u get it?